OpenStudy (anonymous):

when you have sin 2(x/2) is it equivalent to sinx?

OpenStudy (anonymous):

I think yes

OpenStudy (anonymous):

because it is still after sin not before

OpenStudy (anonymous):

Can anyone confirm that it is yes?

OpenStudy (anonymous):

assume that x/2 = 60 degree then sin 2(60) = sin 120 \[\sqrt{3/2}\] and sin 60 = the same value which is \[\sqrt{3/2}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I mean \[\sqrt{3} / 2\]

OpenStudy (anonymous):

No problem your welcome